OpenStudy (anonymous):

3 x + 2 - = ----- 5 6

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

5(x+2)=3(6)

OpenStudy (anonymous):

3/5=x/6+1/3

OpenStudy (anonymous):

x=8/5

OpenStudy (anonymous):

5x+10=18

OpenStudy (anonymous):

5x=8

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

x=8/5

OpenStudy (anonymous):

x=8/5

OpenStudy (anonymous):

or 1.6

OpenStudy (anonymous):

or 1 3/5

OpenStudy (anonymous):

:)